According to an April Associated Press GFK poll, about 30% of Americans support the current health care law. If 10 UNC students were surveyed, find the probability that at most 3 students would support the current health care law. Use 3 decimal places.
Answer #1

p = 0.3, q = 1 - p = 0.7, n = 10

The binomial probability formula is P(x) = C(n, x) p^x q^(n - x)

x P(x) = C(10, x) 0.3^x 0.7^(10 - x)
0 0.0282
1 0.1211
2 0.2335
3 0.2668
4 0.2001
5 0.1029
6 0.0368
7 0.0090
8 0.0014
9 0.0001
10 0.0000

P(At most 3) = P(0) + P(1) + P(2) + P(3) = 0.650